# Penrose’s eight-conic theorem via Penrose’s eight-quadric theorem

### v5

@@ Abstract
This article proves the following theorem, first enunciated by Roger Penrose about 70 years ago: In $\mathbb{R}P^{2}$, if regular conics are assigned to seven of the vertices of a combinatorial cube such that (i) conics connected by an edge are in double contact, and (ii) the chords of contact associated to a cube face meet in a common point, then there exists an eighth conic such that the completed cube satisfies (i) and (ii). This conic is unique if not all the common points of condition (ii) are the same. The proof is based on the following analogous theorem, which is also proved: In $\mathbb{R}P^{3}$, if regular quadrics are assigned to seven of the vertices of a combinatorial cube such that (i) quadrics connected by an edge are in ring contact, and (ii) the ring planes associated to a cube face meet in a common axis, then there exists an eighth quadric such that the completed cube satisfies (i) and (ii). This quadric is unique if not all the common axes of condition (ii) are the same.
@@ 1 Introduction
In 1848, George Salmon published a textbook on conic sections, which includes a further generalization of the Pascal–Brianchon theorem [Sal17, §264], rediscovered by Evelyn, Money-Coutts, and Tyrrell in 1974 [EMCT74, §2.3]. The three point pairs $AA^{\prime}$, $BB^{\prime}$, $CC^{\prime}$ of Pascal’s theorem are degenerations of thin ellipses that are in double contact with the initial conic. And the lines $AB^{\prime}$ and $A^{\prime}B$ are the limit of a pair of common tangents between the two thin ellipses collapsing to $AA^{\prime}$ and $BB^{\prime}$. After replacing the three point pairs with three conics in double contact with a fourth, the remainder of Pappus’s theorem statement still holds (Figure 3).
@@
This article shows a “final” generalization of this progression of projective geometric theorems, first stated by Roger Penrose, the last author of this article, in around 1950 during his undergraduate years. We refer to it Penrose’s eight-conic theorem (Figure 4):
@@
If conics are assigned to seven of the vertices of a combinatorial cube such that conics connected by an edge are in double contact, then (subject to a mild consistency condition on the cube faces) there exists an eighth conic which completes the cube.
@@
That is, if three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ each in double contact with a fourth $\mathbf{{S}}_{0}$, and another three conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ are, respectively, in double contact with $\mathbf{{T}}^{2},\mathbf{{T}}^{3}$, with $\mathbf{{T}}^{3},\mathbf{{T}}^{1}$, and with $\mathbf{{T}}^{1},\mathbf{{T}}^{2}$ (chosen in a consistent branch of a double contact family), then there exists a conic $\mathbf{{T}}^{0}$ that has double contact with $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$.
Here, the consistency condition of each cube face is that the four chords of contact from the four conics have a common point. This condition is automatically guaranteed if one picks a consistent branch of double contact conics with two given conics (out of 3 branches).
This article provides a completion of one of Roger Penrose’s proof ideas for the eight-conic theorem. The idea emerged during his graduate studies at Cambridge that the eight-conic configuration is a planar slice of eight quadrics in ring contact in 3D. (Two quadrics have ring contact if they are tangent to each other along a conic.) It turns out that the following Penrose’s eight-quadric theorem is much more intuitive to see how and why it holds, as shown later in this article:
If quadrics in 3D are assigned to seven of the vertices of a combinatorial cube such that quadrics connected by an edge are in ring contact, then (subject to a mild consistency condition on the cube faces) there exists an eighth quadric which completes the cube.
The face consistency condition is that the four planes containing the four rings of the ring contacts must meet along a common axis. One advantage of the 3D approach is that this face consistency condition presents a much less serious obstacle to completing the cube than the 2D face consistency condition.
The article is structured as follows. Section 2 provides the necessary background information on conics and quadrics. In Section 3, we present the main theorems: the eight-conic theorem and its 3D counterpart, the eight-quadric theorem. The proof of the eight-quadric theorem is given in Section 3.4, with the eight-conic theorem derived as a corollary through an extrusion process explained in Section 3.4.4. Section 3.5 illustrates examples of the eight-quadric configuration, showing how its 2D slices or projections yield classical theorems in the projective plane. Section 4 presents parametrized formulas for the conics of the combinatorial cube in terms of a minimal set of initial data. Section 5 of the article provides a brief account of related research and of the genesis of this article. Section 6 considers possible directions for further research.
@@ 2.1.1 Pencils of conics
The linear span of two conics $A$ and $B$ is called a pencil of conics. A pencil that contains a regular conic is called a regular pencil. We will need the following lemma in Sect. 3.2:
Lemma 2.1 (Regular pencil).
@@ 2.3 Notation
Before we turn to the proof of the theorems, we introduce some notation used in the sequel. First, an alternative notation for degenerate conics, used in Sec. 3.2 and Sec. 4. Secondly, notation for so-called complete quadrics, required to represent non-regular conics and quadrics, and used in Sec. 3.3.
@@ 2.3.1 Abbreviated notation
Until now we have written a line pair formed by the two lines $p$ and $q$ directly as a matrix via the symmetrized outer project $pq^{T}+qp^{T}$, and the double line of $p$ as $pp^{T}$. $p$ and $q$ can also be interpreted as the zero sets of the linear polynomials
@@
The final equality shows that $r(x,y,z)=0$ represents the same line pair as $(pq^{T}+qp^{T})$. Setting $q=p$ yields an expression for the double line $p^{2}$ which agrees with the matrix $pp^{T}$. We introduce the expressions $\mathbf{pq}$ and $\mathbf{p}^{2}$ as the abbreviated notation for the line pair, resp., the double line. We also write $\mathbf{p}$ for the line itself (not a double line); although it is not a conic, it can be used to generate conics within the ring of polynomials in the variables $\{x,y,z\}$. We continue to use $p$ to denote a line in $\mathbb{R}P^{2}$ when the membership in the polynomial ring is not highlighted.
@@ 2.3.2 Notation for complete quadrics
Every quadric has a point-wise, or primary, aspect as well as a plane-wise, or dual, aspect. We write the primary (dual) form of a quadric as $\mathbf{{Q}}$ ($\mathbf{{Q}}^{*}$). When does the pair $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ represent a quadric? When either is regular, then so is the other, and the relationship is given by $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda\neq 0$. Otherwise, both are degenerate and satisfy the relationship $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda=0$, which can be seen by taking limits of regular quadrics. Then a necessary condition for $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ to represent a quadric is $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$. If in addition both are not zero, we call the pair a complete quadric. In discussions limited to regular quadrics, we usually refer to a quadric simply in its point-wise form as $\mathbf{{Q}}$, since $\mathbf{{Q}}^{*}$ is uniquely determined. When degenerate quadrics play a role, as in Sec. 3.3, we employ the full notation introduced here.
@@ 3 The main theorems
Theorem 3.1 (8-conic theorem).
In $\mathbb{R}P^{2}$, if regular conics are assigned to seven of the vertices of a combinatorial cube such that (i) conics connected by an edge are in double contact, and (ii) the chords of contact associated to a cube face meet in a common point, then there exists an eighth conic such that the completed cube satisfies (i) and (ii). This conic is unique if not all the common points of condition (ii) are the same.
Theorem 3.2 (8-quadric theorem).
In $\mathbb{R}P^{3}$, if regular quadrics are assigned to seven of the vertices of a combinatorial cube such that (i) quadrics connected by an edge are in ring contact, and (ii) the ring planes associated to a cube face meet in a common axis, then there exists an eighth quadric such that the completed cube satisfies (i) and (ii). This quadric is unique if not all the common axes of condition (ii) are the same.
@@ 3.1 Notation and terminology
Both the 8-conic and 8-quadric theorems assume that a regular Penrose 7-configuration is given and assert a unique completion to an 8-configuration. Note that this eighth conic (or quadric) is not necessarily unique or regular (see Thm. 3.3).
@@
Next, we introduce notation for the vertices and faces of the cube. Consult Fig. 5, left. We decompose the cube vertices into two interlocked tetrahedra and label the first group of vertices $\mathbf{{S}}_{i},i\in(0,1,2,3)$ and the second $\mathbf{{T}}^{i}$, where $\mathbf{{T}}^{i}$ is opposite to $\mathbf{{S}}_{i}$ with respect to the center of the cube. Write the face of the cube with corners $\mathbf{{T}}^{i}\mathbf{{S}}_{k}\mathbf{{T}}^{j}\mathbf{{S}}_{m}$ as $F^{ij}_{km}$ where $ijkm$ is a permutation of $0123$.
In condition (ii) of the 8-conic theorem, we call the common point of the chords of contact on a face the face point of the face; for condition (ii) of the 8-quadric theorem, we call the common axis of the ring planes the face axis of the face.
Furthermore, in what follows, we assume that the empty vertex of a Penrose 7-configuration is $\mathbf{{T}}^{0}$.
@@ 3.2 The face figure of a Penrose conic configuration
In Fig. 5 (left), consider two conics corresponding to opposite nodes in a face of the cube, for example, $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$. Then by (6) they can be written as $\mathbf{{T}}^{1}=\mathbf{{S}}_{0}+\mathbf{q}^{2}$ and $\mathbf{{T}}^{3}=\mathbf{{S}}_{0}+\mathbf{q}^{2}$ with $\mathbf{p}$ and $\mathbf{q}$ being the chords of contact between $\mathbf{{S}}_{0}$ and $\mathbf{{T}}^{i}$. Using abbreviated notation:
@@
where $\mathbf{d}:=\mathbf{p}+\mathbf{q}$ and $\mathbf{e}=\mathbf{p}-\mathbf{q}$ is a line pair in the pencil spanned by $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$. Hence $\mathbf{p},\mathbf{q},\mathbf{d}$ and $\mathbf{e}$ have a common point. Consult Fig. 5, right.
By Lemma 2.1, there are at most 3 line pairs in a regular pencil of conics. Hence, the set of conics that are in double contact with $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$ splits into at most 3 families such that condition (ii) of the 8-conic theorem holds automatically for all conics $\mathbf{{S}}_{0},\mathbf{{S}}_{2}$ in the same family. All these conics can be explicitly written in terms of $\mathbf{p}$, $\mathbf{q}$, and $\mathbf{{S}}_{0}$ (see [Sal17, §287]).
@@ 3.3 Penrose’s 8-conic theorem
The extrusion needed for the 7-configurations encountered in the following proofs is provided by the following lemma, whose proof appears below in Sect. 3.4.4.
Lemma 3.1 (Extrusion).
@@
Apply Lemma 3.1 to the given regular Penrose 7-conic configuration to obtain an extrusion to a regular Penrose 7-quadric configuration. Then apply Thm. 3.2 to complete this to a Penrose 8-quadric configuration. The intersection of the plane of the original 7-conic configuration with the complete 8-quadric configuration is the desired regular Penrose 8-conic configuration. Face points of the conic configuration are distinct exactly when the corresponding face axes of the quadric configuration are distinct. Hence the eighth conic is unique exactly when the face points are not all the same. ∎
@@ 3.4 Penrose’s 8-quadric theorem
Let $i$ and $j\in\{0,1,2,3\}$, $m$ and $n$ in $\{1,2,3\}$ and $P^{m}_{i}$ denote the ring point for $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{m}$, $p^{m}_{i}$ denote the corresponding ring plane, $\Lambda^{mn}_{ij}$ the common axis of the ring planes for the face $F^{mn}_{ij}$ and $\lambda^{mn}_{ij}$ the common spear (point-wise line) of the ring points. We call these the face axes, resp., the face spears. By assumption, they are well defined. Equivalently, given any one of the three complete faces of the 7-configuration, its four ring planes are not all identical. Consider, for instance, the face $F^{12}_{03}$. By Lemma 2.4, we immediately conclude that $\mathbf{{S}}_{0}$ is not in ring contact with $\mathbf{{S}}_{3}$ and $\mathbf{{T}}^{1}$ is not in ring contact with $\mathbf{{T}}^{2}$. By similar reasoning applied to the other two complete faces, we get the following lemma:
Lemma 3.2.
@@
These will both be non-zero since otherwise, if we had $P^{n}_{i}=P^{n}_{j}$ or $p^{n}_{i}=p^{n}_{j}$, $\mathbf{{S}}_{i}$ and $\mathbf{{S}}_{j}$ would have to be in ring contact along the common conic of intersection with $\mathbf{{T}}^{n}$, contradicting Lemma 3.2.
@@
Lemma 3.3 (Concurrent axes and spears).
@@
We show now that the $\Lambda^{ij}_{0k}$ cannot all lie in a plane. Indeed, if they do, then that plane must be identical to the three ring planes $p^{j}_{0}$ but by the regularity of $\mathbf{{S}}_{0}$ this would imply that $\mathbf{{T}}^{i}$ is in ring contact to $\mathbf{{T}}^{j}$ for $i,j=1,2,3$ contradicting Lemma 3.2.
@@
By regularity and duality the analogous result can be obtained for the ring points and the face spears. Furthermore, by Lemma 2.3, $O$ (resp., $\Omega$) and $o$ (resp., $\omega$) are a polar pair with respect to all quadrics of the configuration.
@@
Lemma 3.4 (Conical Octahedron Lemma).

### v7

@@ Abstract
This article proves the following theorem, first enunciated by Roger Penrose about 70 years ago but never published: In $\mathbb{R}P^{2}$, if conics are assigned to seven of the vertices of a combinatorial cube such that (i) conics connected by an edge are in double contact, and (ii) the chords of contact associated to a cube face meet in a common point, then there exists an eighth conic such that the completed cube satisfies (i) and (ii). The theorem turns out to be a remarkable generalization of many well-known theorems of projective geometry – Pappus, Desargues, Pascal, Brianchon and Monge are the best-known ones. This archetypal principle provides a unifying framework in which the myriad specializations of the theorem and their interrelationships can be grasped as an organic whole, enriching the field of projective geometry and opening new vistas for research. The article begins with a series of motivational examples. It then gives a geometric proof assuming that the conics are regular, followed by an algebraic one that removes this restriction. The geometric proof is obtained as a slice of an analogous theorem for quadrics in $\mathbb{R}P^{3}$; the algebraic one is based on the determinants of a special matrix associated to the configuration of conics.
@@ 1 Introduction
In 1848, George Salmon published a textbook on conic sections, which includes a further generalization of the Pascal–Brianchon theorem [Sal17, §264], rediscovered by Evelyn, Money-Coutts, and Tyrrell in 1974 [EMCT74, §2.3]. The three point pairs $AA^{\prime}$, $BB^{\prime}$, $CC^{\prime}$ of Pascal’s theorem are degenerations of thin ellipses that are in double contact with the initial conic. And the lines $AB^{\prime}$ and $A^{\prime}B$ are the limit of a pair of common tangents between the two thin ellipses collapsing to $AA^{\prime}$ and $BB^{\prime}$. After replacing the three point pairs with three conics in double contact with a fourth, the remainder of Pappus’s theorem still holds (Figure 3).
@@ 2 Overview
This article presents a “final” generalization of this progression of projective geometric theorems, first stated by Roger Penrose, the last author of this article, in around 1950 during his undergraduate years. We refer to it Penrose’s eight-conic theorem (Figure 4, right):
@@
If conics are assigned to seven of the vertices of a combinatorial cube such that conics connected by an edge are in double contact, and the chords of contact in each face share a common point, then there exists an eighth conic which completes the cube under the same conditions.
@@
That is, if three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ each in double contact with a fourth $\mathbf{{S}}_{0}$, and another three conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ are, respectively, in double contact with $\mathbf{{T}}^{2},\mathbf{{T}}^{3}$, with $\mathbf{{T}}^{3},\mathbf{{T}}^{1}$, and with $\mathbf{{T}}^{1},\mathbf{{T}}^{2}$ (so that the chords of contact of a face share a point), then there exists a conic $\mathbf{{T}}^{0}$ that has double contact with $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$. We will show below in Section 4.2 that the condition on the chords of contact is equivalent to choosing the correct branch among 3 possible families of double contact conics.
The article first presents a geometric proof of the theorem, under the assumption that the conics are regular. After reviewing background material on projective geometry (Section 3), it proceeds by proving an analogous “8-quadric” theorem, in which conics in double contact are replaced with 3D quadrics in ring contact (Section 4). The 8-conic theorem is proved from the 8-quadric theorem by showing that 1) the slice of a regular 8-conic configuration with a generic plane is a regular 8-conic configuration, and 2) a regular 7-conic configuration can be extruded to a regular 7-quadric configuration. Proving the regular 8-conic theorem in this way has the advantage that in 3D there is extra geometric structure that simplifies the proof. The idea for this proof originated with Roger Penrose during his graduate studies and was reconstructed for this article.
The geometric proof is followed by a discussion of special 3D configurations that yield, by slicing, well-known theorems in the projective plane (Section 5), introducing the possibility of non-regular conics in the cube.
Equipped with a geometric understanding of the content of the 8-conic theorem in this weaker form, the article proceeds in Section 6 to an algebraic proof that applies to all conics, regular or not; for example, the line pairs and point pairs of the example theorems above, or even double lines and double points. The proof is carried out on the basis of a series of parametrized formulas relating the neighboring conics of the cube. Section 7 shows that all these formulas can be obtained as subdeterminants of a special $4\times 4$ symmetric matrix, some of whose entries are linear or quadratic polynomials. The symmetric subdeterminants correspond to the conics of the Penrose cube, while the other subdeterminants contain accessory information, such as the chords of contact, or even secondary theorems not contained in the 8-conic theorem itself.
Section 8 of the article gives a brief account of previous research and of the genesis of this article. Finally, Section 9 provides a recapitulation of the article integrated with a discussion of directions for future research.
@@ 3.1.1 Pencils of conics
The linear span of two conics $A$ and $B$ is called a pencil of conics. A pencil that contains a regular conic is called a regular pencil. We will need the following lemma in Sect. 4.2:
Lemma 1 (Regular pencil).
@@ 3.3 Notation
Before we turn to the proof of the theorems, we introduce some notation used in the sequel. First, an alternative notation for degenerate conics, used in Sec. 4.2 and Sec. 6. Secondly, notation for so-called complete quadrics, required to represent non-regular conics and quadrics, and used in Sec. 14.
@@ 3.3.1 Abbreviated notation
Until now, we have written a line pair formed by the two lines $p$ and $q$ directly as a matrix via the symmetrized outer project $pq^{T}+qp^{T}$, and the double line of $p$ as $pp^{T}$. $p$ and $q$ can also be interpreted as the zero sets of the linear polynomials
@@
The final equality shows that $r(x,y,z)=0$ represents the same line pair as $(pq^{T}+qp^{T})$. Setting $q=p$ yields an expression for the double line $p^{2}$ which agrees with the matrix $pp^{T}$. We introduce the expressions $\mathbf{pq}$ and $\mathbf{p}^{2}$ as the abbreviated notation for the line pair, resp., the double line. We also write $\mathbf{p}$ for the line itself (not a double line); although it is not a conic, it will be used in Section 7 to generate conics within the ring of polynomials in the variables $\{x,y,z\}$. We continue to use $p$ to denote a line in $\mathbb{R}P^{2}$ when the membership in the polynomial ring is not highlighted.
@@ 3.3.2 Notation for complete quadrics
Every quadric has a point-wise, or primary, aspect as well as a plane-wise, or dual, aspect. We write the primary (dual) form of a quadric as $\mathbf{{Q}}$ ($\mathbf{{Q}}^{*}$). When does the pair $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ represent a quadric? When either is regular, then so is the other, and the relationship is given by $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda\neq 0$. Otherwise, both are degenerate and satisfy the relationship $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$ for $\lambda=0$, which can be seen by taking limits of regular quadrics. Then a necessary condition for $(\mathbf{{Q}},\mathbf{{Q}}^{*})$ to represent a quadric is $\mathbf{{Q}}\mathbf{{Q}}^{*}=\lambda Id$. If in addition both are not zero, we call the pair a complete quadric. In discussions limited to regular quadrics, we usually refer to a quadric simply in its point-wise form as $\mathbf{{Q}}$, since $\mathbf{{Q}}^{*}$ is uniquely determined. When degenerate quadrics play a role, as in Sec. 14, we employ the full notation introduced here.
@@ 4 The main theorems
We state the 8-conic and 8-quadric theorems in their full generality now.
Theorem 7 (8-conic theorem).
In $\mathbb{R}P^{2}$, if conics are assigned to seven of the vertices of a combinatorial cube such that (i) conics connected by an edge are in double contact, and (ii) the chords of contact associated to a cube face meet in a common point, then there exists an eighth conic such that the completed cube satisfies (i) and (ii). This conic is unique if not all the common points of condition (ii) are the same.
Theorem 8 (8-quadric theorem).
In $\mathbb{R}P^{3}$, if quadrics are assigned to seven of the vertices of a combinatorial cube such that (i) quadrics connected by an edge are in ring contact, and (ii) the ring planes associated to a cube face meet in a common axis, then there exists an eighth quadric such that the completed cube satisfies (i) and (ii). This quadric is unique if not all the common axes of condition (ii) are the same.
We begin with a geometric proof under the weaker condition that the conics are regular. This will be followed in Section 6 with an algebraic proof that removes this restriction. We first prove Thm. 8 and use it to prove Thm 7.
@@ 4.1 Notation and terminology
Both the 8-conic and 8-quadric theorems assume that a Penrose 7-configuration is given and assert a unique completion to an 8-configuration.
@@
Next, we introduce notation for the vertices and faces of the cube. Consult Figure 5, left. We decompose the cube vertices into two interlocked tetrahedra and label the first group of vertices $\mathbf{{S}}_{i},i\in(0,1,2,3)$ and the second $\mathbf{{T}}^{i}$, where $\mathbf{{T}}^{i}$ is opposite to $\mathbf{{S}}_{i}$ with respect to the center of the cube. Write the face of the cube with corners $\mathbf{{T}}^{i}\mathbf{{S}}_{k}\mathbf{{T}}^{j}\mathbf{{S}}_{m}$ as $\mathcal{F}^{ij}_{km}$ where $ijkm$ is a permutation of $0123$. The common point of the chords of contact in this face we write as $P^{ij}_{km}$, the face point of the face. Since the edges of the cube are double contact pencils, each edge contains a unique double line $\mathbf{{V}}^{i}_{j}$, the chord of contact of $\mathbf{{T}}^{i}$ and $\mathbf{{S}}_{j}$.
When working with the 8-quadric theorem, we call the common axis of the ring planes of the face $\mathcal{F}^{ij}_{km}$ the face axis $\Lambda^{ij}_{km}$. Furthermore, in what follows, we assume that the empty vertex of a Penrose 7-configuration is $\mathbf{{T}}^{0}$.
@@ 4.2 The face figure of a Penrose conic configuration
In Figure 5 (left), consider two conics corresponding to opposite nodes in a face of the cube, for example, $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$. Then by (6) they can be written as $\mathbf{{T}}^{1}=\mathbf{{S}}_{0}+\mathbf{p}^{2}$ and $\mathbf{{T}}^{3}=\mathbf{{S}}_{0}+\mathbf{q}^{2}$ with $\mathbf{p}$ and $\mathbf{q}$ being the chords of contact between $\mathbf{{S}}_{0}$ and $\mathbf{{T}}^{i}$. Using abbreviated notation:
@@
where $\mathbf{d}:=\mathbf{p}+\mathbf{q}$ and $\mathbf{e}=\mathbf{p}-\mathbf{q}$ is a line pair in the pencil spanned by $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$. Hence $\mathbf{p},\mathbf{q},\mathbf{d}$ and $\mathbf{e}$ have a common point. Consult Figure 5, right.
By Lemma 1, there are at most 3 line pairs in a regular pencil of conics. Hence, the set of conics that are in double contact with $\mathbf{{T}}^{1}$ and $\mathbf{{T}}^{3}$ splits into at most 3 1-parameter families such that condition (ii) of the 8-conic theorem holds automatically for all conics $\mathbf{{S}}_{0},\mathbf{{S}}_{2}$ in the same family. Each 1-parameter family can be explicitly written in terms of $\mathbf{p}$, $\mathbf{q}$, and $\mathbf{{S}}_{0}$ (see [Sal17, §287]). Given $\mathbf{{S}}_{1}$ in such a family, its parameter value can be obtained directly from this explicit form.
@@ 4.3 Penrose’s 8-conic theorem
The extrusion needed for the 7-configurations encountered in the following proofs is provided by the following lemma, whose proof appears below in Sect. 4.4.4.
Lemma 10 (Extrusion).
@@
Apply Lemma 10 to the given regular Penrose 7-conic configuration to obtain an extrusion to a regular Penrose 7-quadric configuration. Then apply Thm. 8 to complete this to a Penrose 8-quadric configuration. The intersection of the plane of the original 7-conic configuration with the complete 8-quadric configuration is the desired regular Penrose 8-conic configuration. Face points of the conic configuration are distinct exactly when the corresponding face axes of the quadric configuration are distinct. Hence the eighth conic is unique exactly when the face points are not all the same. ∎
@@ 4.4 Penrose’s 8-quadric theorem
Let $i$ and $j\in\{0,1,2,3\}$, $m$ and $n$ in $\{1,2,3\}$ and $P^{m}_{i}$ denote the ring point for $\mathbf{{S}}_{i}$ and $\mathbf{{T}}^{m}$, $p^{m}_{i}$ denote the corresponding ring plane, $\Lambda^{mn}_{ij}$ the common axis of the ring planes for the face $F^{mn}_{ij}$ and $\lambda^{mn}_{ij}$ the common spear (point-wise line) of the ring points. We call these the face axes, resp., the face spears. By assumption, they are well defined. Equivalently, given any one of the three complete faces of the 7-configuration, its four ring planes are not all identical. Consider, for instance, the face $\mathcal{F}^{12}_{03}$. By Lemma 5, we immediately conclude that $\mathbf{{S}}_{0}$ is not in ring contact with $\mathbf{{S}}_{3}$ and $\mathbf{{T}}^{1}$ is not in ring contact with $\mathbf{{T}}^{2}$. By similar reasoning applied to the other two complete faces, we get the following lemma:
Lemma 11.
@@
These will both be non-zero since otherwise, if we had $P^{n}_{i}=P^{n}_{j}$ or $p^{n}_{i}=p^{n}_{j}$, $\mathbf{{S}}_{i}$ and $\mathbf{{S}}_{j}$ would have to be in ring contact along the common conic of intersection with $\mathbf{{T}}^{n}$, contradicting Lemma 11.
@@
Lemma 12 (Concurrent axes and spears).
@@
We show now that the $\Lambda^{ij}_{0k}$ cannot all lie in a plane. Indeed, if they do, then that plane must be identical to the three ring planes $p^{j}_{0}$ but by the regularity of $\mathbf{{S}}_{0}$ this would imply that $\mathbf{{T}}^{i}$ is in ring contact to $\mathbf{{T}}^{j}$ for $i,j=1,2,3$ contradicting Lemma 11.
@@
By regularity and duality the analogous result can be obtained for the ring points and the face spears. Furthermore, by Lemma 3, $O$ (resp., $\Omega$) and $o$ (resp., $\omega$) are a polar pair with respect to all quadrics of the configuration.
@@
Lemma 13 (Conical Octahedron Lemma).